OpenStudy (anonymous):

Someone please halp! what are the x-intercepts of the graph of y+12=x^2+x

OpenStudy (anonymous):

first of all you want to isolate y. y = x^2 + x - 12 Now you can use the quadratic formula to solve it. x^2 + x - 12 = 0 Do you know the formula?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

Maybe there is a better solution for this particular case but the general solution is \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] Do you recognize this?

OpenStudy (anonymous):

oh ok yes i do

OpenStudy (anonymous):

Good, try to plug it in and see what you get for x :)

OpenStudy (anonymous):

ok thanks so much!

OpenStudy (precal):

you can also, factor it as well